# A Comprehensive Multi-Vocal Empirical Study of ML Cloud Service Misuses

### v1

@@ 3.1. Systematic Literature Review
Second, we performed this search query on popular digital libraries for research papers, including ACM, IEEE Xplore, and Google Scholar, leading to an initial list of 9,054 papers. The search has been refined by applying filters based on relevance and a publication date range limited to the past ten years. We then filtered these collected papers according to their titles and abstracts. Two of the coauthors manually and independently analyzed all the papers and then reconciled any differences through discussion meetings. We included in our review the papers that met at least two of the following criteria: (1) papers related to ML; (2) papers related to cloud computing services; (3) papers related to (anti)patterns. Finally, we applied forward and backward snowballing techniques (wohlin2014guidelines, ) to minimize the risk of missing relevant papers. Our final set comprises a collection of 39 research papers published between 2017 and 2024. The distribution of these papers across years is depicted in Figure 2. Notably, the highest number of ML service misuses reported in a single study within the existing literature is eight, as provided by Wan et al. (wan2021machine, ). This highlights the need for a more comprehensive and larger catalog of ML misuses, addressing the identified research gap in prior work, as we propose in our study.
Figure 2. Number of Research Papers Published Per Year
3.2. Gray Literature Review
In addition to our systematic literature review, we conducted an in-depth examination of the official documentation provided by major ML cloud providers, including Microsoft, Amazon, and Google. We compiled a detailed list of the ML services offered by each provider and carefully analyzed their documentation. Our focus was twofold: identifying recommended best practices and uncovering non-recommended practices that could affect the performance of ML service-based systems in terms of model accuracy, cloud resource usage, and latency.
To gain deeper insights into the challenges and misuses developers encounter when working with ML services, we also reviewed a wide range of online resources. These included online blogs from the Microsoft Tech Community  [^1] https://techcommunity.microsoft.com, Google Developers [^2] https://developers.googleblog.com/en, and AWS Developer platforms  [^3] https://aws.amazon.com/blogs/developer, as well as publications on Medium and discussions on Stack Overflow. We selected these platforms due to their widespread popularity and extensive repositories of blogs and discussions that frequently address programming challenges in general, and ML-related issues in particular.
We should note that the systematic literature review and the gray literature review provided valuable insights into the practical use of ML services within software systems. By integrating these data, we were able to derive a first set of ML service misuses, which formed a preliminary catalog of 37 misuses.
3.3. Review of Open-Source ML Cloud Service-Based Systems
@@
To collect concrete examples of ML service misuses, we conducted an in-depth analysis of various GitHub repositories. We first started with the dataset provided by Wan et al. (wan2021machine, ) and observed that it contains 307 Python projects covering eight ML service misuses. Moreover, their dataset focused exclusively on ML APIs from Google and Amazon, which limited the scope of ML applications. Given the unavailability of datasets that meet our needs, we proceeded to collect additional GitHub projects, as described below.
@@
    Projects Collection: We collected ML service-based projects on GitHub, a rich source of open-source ML-based software systems. The data collection and analysis was from May 2024 to July 2024, without a specific time frame for the creation date of projects. We used the GitHub API  [^4] https://docs.github.com/en/rest to automatically retrieve projects, enabling a more efficient and structured retrieval of information. We used search keywords listed by cloud service providers. In addition, we employed GitHub’s search functionality to identify the presence of those keywords within Python files. For example, to collect repositories using Microsoft Azure ML, we considered the following keywords: “ML cloud”, “Azure cognitive service”, “API”, “Azure AI”, and “AutoML”. The collected information about the projects includes the repositories’ URLs, number of stars, number of forks, contributors, and descriptions.
@@
    Projects Filtration: We filtered the collected repositories by automatically analyzing the project description and source code. This involved inspecting the development language, identifying the presence of Jupyter notebooks, and locating Python files. This allowed us to filter repositories based on their relevance to our study. The data collection process resulted in 1,900 Python GitHub repositories related to ML services. We randomly selected a sample of these projects (377) for further manual analyses, as we aimed to cover repositories related to major cloud providers (e.g., Microsoft, Google, and Amazon) and ML services (e.g., text-to-speech, and image recognition) to ensure a broad representation. However, we found that not all repositories mentioning cloud services or ML-related terms really use ML services. To maintain the focus on projects that use cloud-based ML services, we excluded repositories that met at least one of the following criteria:
    • Toy projects presented as small demonstrations or tutorials.
    • Official Toolkits from Cloud Providers: These repositories often serve as templates or tools rather than complete projects with fully implemented use cases.
    • Misleading Mention: Projects that mention cloud providers or ML-related terms but do not concretely use any ML service, such as Vertex AI, AWS Sagemaker, or Azure ML, in their source code. For example, we excluded a repository deploying a microservice on Google Cloud but not involving any ML service, despite mentioning Google AI in the ReadMe file.
    Applying the exclusion criteria was done manually and resulted in a curated set of 87 ML service-based projects for further analysis.
@@
    Projects Analyses: Subsequently, two of the coauthors (with experience in machine learning and software engineering) played the role of evaluators and manually and independently analyzed the 87 projects to detect the presence of different ML service misuses. We used Understand [^5] https://scitools.com, a static code analysis tool, to perform a thorough code review and control flow graph generation in complex repositories containing numerous directories. This tool provides different views of the architecture, which enabled a comprehensive understanding and analysis of the system.
@@
    Misuses Isolation: As part of the detection process, the two coauthors independently and systematically identified and highlighted instances of ML service misuse, categorizing them by their stage of ML development pipeline. Each detected misuse is accompanied by a detailed explanation, which provides reasons for why the identified instance is problematic and a link to the exact line of code within the repository, facilitating quick access for further inspection. For example, to detect “Non specification of early stopping criteria” misuse, the two coauthors independently analyzed the ML service-based system, identified the ML service cloud provider used, verified the training component of the code, and reviewed the recommendations for specifying early stopping criteria based on the official documentation of the cloud provider. If the training component did not adhere to best practices for cost-efficiency and resource optimization, a misuse was detected.
@@
    Detection Agreement: The two coauthors tried to resolve any disagreement during the process through discussion meetings. In cases where a consensus could not be reached, a third coauthor was involved to provide additional insights and resolve disagreements. To assess the level of agreement between the evaluators, we calculated Cohen’s kappa coefficient (banerjee1999beyond, ), a statistic used to measure inter-rater reliability. We achieved Cohen’s kappa of 84.7%, indicating a high level of agreement among the evaluators and highlighting the reliability of the manual identification process of the occurrences of ML service misuses.
  6. f)
    Detection Results: Our manual evaluation of the projects on GitHub resulted in the identification of 20 misuses (with 290 occurrences) of ML service misuses in our dataset. To the best of our knowledge, this represents the largest catalog of ML service-based misuses based on various source: systematic and gray literature as well as open-source project analysis (involving both automated and manual). This sets a new benchmark for research in this area. Figure 3 illustrates the prevalence of these misuses, along with the distribution of their occurrences.
Figure 3. Occurrences of ML service misuses in our studied projects, including the misuses identified in the systematic literature [SL] and the gray literature [GL]
3.4. Pre-validation of the Catalog of ML Service Misuses
We curated a preliminary catalog of a total of 34 ML service misuses: 14 misuses were identified through our systematic review of the research literature and 25 through our examination of the gray literature, with an overlap of five misuses. Then, our analysis of GitHub projects uncovered 20 misuses, which included six from the research literature and 14 from the gray literature.
@@ 3.4. Pre-validation of the Catalog of ML Service Misuses
The pre-validation of the catalog involved multiple iterations. The two coauthors independently reviewed and evaluated the relevance of the identified misuses. They applied the following inclusion criteria to retain: (1) misuses specific to ML service-based systems, excluding those applicable to general service systems; (2) misuses applicable to all cloud providers; (3) misuses with occurrences identified within GitHub projects; and (4) misuses that are potentially detectable through source code analysis. They also classified the misuses according to their stage within the ML development pipeline. After such validation, we retained 20 ML service misuses, which we will describe in detail in Section 4.1.
3.5. Survey
We conducted an online survey with practitioners to validate our catalog and get better insight into the usage of ML cloud services in industry. The survey consisted of two main phases:
@@ 3.5. Survey
Preparation of the Online Survey: We created a web-based survey (googleform2025, ) using Google Forms. We designed our survey questions based on the systematic literature review, the gray literature review, and the analysis of ML service-based systems. This helped us identify and shape the different parts of the survey, their various questions, and the response options for each question. Before publishing the survey, we conducted a pilot interview with three professionals and four researchers to identify any duplicated or inconsistent questions in the survey before distribution. This also helped to assess the survey length, which might discourage participants from responding, and to estimate the approximate time required to complete the survey. The seven participants went through the survey and suggested minor modifications. To facilitate participants’ understanding of ML service misuses, we included a link to our online catalog [^6] https://ml-service-misuses.github.io/Catalog-Proposal in the survey to further explain each misuse. The final survey consisted of seven sections, primarily about the demographics of the participants, the frequency of usage of ML services, the introduction of ML misuses, the agreement level with the catalog, and the possible reasons for misuses.
@@
Selection of Participants: We targeted practitioners with experience in ML cloud services. Identifying and inviting such practitioners was challenging due to the key requirement for familiarity with both ML and cloud computing, as well as practical experience with ML services in their work. To select the participants, we relied on (1) information about companies that offer ML cloud services and their clients/partners, (2) the authors’ networks (e.g., former colleagues), and (3) search queries on LinkedIn profiles (e.g., AI Engineer OR Cloud Architect OR Azure certified OR Cloud for AI). For those who were identified as potential practitioners, we sent them invitations via E-Mail, LinkedIn, and Facebook. We chose to limit our invitations to a maximum of five practitioners per organization to ensure a wide representation and mitigate the risk of bias. We invited 421 practitioners to complete the survey and encouraged them to share our invitations within their network. The survey was completed by 50 practitioners, which corresponds to a completion rate of 12%.
Statistical Correlation: We measured the Spearman’s correlation (ali2022spearman, ) between (a) participants’ experience levels in AI development and the number of ML service misuses they agreed upon, (b) participants’ experience levels and the frequency of encountering each ML service misuse. We used the Spearman’s correlation as it does not require assumptions about data distribution.

### v2

@@ 4.1. Research Literature Review (RQ1)
We executed this query on popular digital libraries for research papers, including ACM, IEEE Xplore, Scopus, and Springer, leading to an initial list of 786 papers. The search has been refined by applying filters based on relevance and a publication date range limited to the past ten years. We then filtered these collected papers according to their titles and abstracts. Two of the co-authors (with experience in cloud-based ML services research) independently conducted a manual review of the papers, resolving disagreements through discussion to reach consensus. We calculated the Cohen’s kappa coefficient (banerjee1999beyond, ), a statistic commonly used to measure inter-rater reliability. The resulting value was 89.79%, indicating a strong level of agreement between the reviewers. We included in our review those papers that met at least two of the following criteria: (1) studies related to ML; (2) studies related to cloud computing services; and (3) studies related to (anti)patterns. Finally, to minimize the risk of missing relevant studies, we applied forward and backward snowballing techniques (wohlin2014guidelines, ), conducting eight iterations that resulted in the inclusion of six additional studies.
4.2. Gray Literature Review (RQ1)
In addition to our research literature review, we conducted a review of the gray literature on ML cloud service misuses, following the guidelines outlined by Garousi et al. (garousi2019guidelines, ). We started by searching for arXiv papers using the same search query terms and inclusion criteria as in our research literature review to ensure consistency in our scope. We then expanded our search to include official documentation and practical developer insights from various online sources, executing the same query on Google Search to capture a broad spectrum of gray literature.
Our search query led to the identification of 3,116 online references. To select only relevant resources, we applied the following selection criteria:
  • (1)
    We filtered out non-English content to ensure clarity and consistency.
  • (2)
    We selected resources from only highly credible platforms known for their rich, practical insights and widespread developer usage. These platforms include (1) Official Documentation from major ML cloud providers: Microsoft, Amazon, and Google; and (2) Official Cloud Technical Blogs and Developer Platforms: Microsoft Tech Community [^1] https://techcommunity.microsoft.com, Google Developers Blog [^2] https://developers.googleblog.com/en, AWS Developer Blogs [^3] https://aws.amazon.com/blogs/developer, Medium Blogs [^4] https://medium.com; and (3) Specialized books that cover the design and usage of ML cloud services (tranquillin2023architecting, ; ping2022machine, ). These sources provide diverse perspectives, ranging from official best-practice recommendations to real-world challenges and common misuses reported by developers.
  • (3)
    We selected online references that were relevant to the context of our study and aligned with the latest developments in ML cloud services.
  • (4)
    We removed duplicates and outdated materials, specifically those published before the most recent stable release of each cloud or ML service at the time of our research, to ensure up-to-date, quality sources.
We performed five iterations of backward and forward snowballing, which yielded 12 additional relevant sources on ML cloud service misuses. We should note that to ensure consistency with our research literature review, we applied a publication date filter from 2017 to 2024, as this period reflects the most relevant and recent retained papers. Moreover, to ensure a focused and manageable review, we reviewed references until we found mostly non-relevant links. Specifically, after reviewing the first 10% of references in each source and encountering ten consecutive non-relevant links, we stopped further analysis. This approach allowed us to retain the most relevant resources while keeping the review practically feasible and representative.
4.3. Analysis of GitHub Projects (RQ1 & RQ3)
@@
To collect concrete examples of ML service misuses (RQ1) and study their prevalence (RQ3) in open-source software, we conducted an in-depth analysis of various repositories hosted on GitHub, while relying on the preliminary catalog of the misuses identified in the previous steps. We initially started with the dataset provided by Wan et al. (wan2021machine, ) and observed that it contains 307 Python projects covering eight ML service misuses. Moreover, their dataset focused exclusively on ML APIs from Google and Amazon, which limited the scope of ML applications. Considering the lack of similar datasets that satisfy our needs, we proceeded to collect additional GitHub projects, as described below.
@@
    Projects Collection: We collected ML service-based GitHub projects, a rich source of open-source ML-based software systems. The data collection and analysis were carried out during the period between May 2024 and July 2024, without a specific time frame for the creation date of such projects. We used the GitHub API [^5] https://docs.GitHub.com/en/rest to automatically retrieve projects, enabling a more efficient and structured retrieval of information. We used specific search keywords listed by the cloud service providers under study. In addition, we employed GitHub’s search functionality to identify the presence of those keywords within Python files. For example, to collect repositories using Microsoft Azure ML, we considered the following keywords: “ML cloud”, “Azure cognitive service”, “API”, “Azure AI”, and “AutoML”. The collected information about the projects includes the repositories’ URLs, number of stars, number of forks, contributors, and descriptions.
@@
    Projects Filtration: We filtered the collected repositories by automatically analyzing the project description and source code. This involved inspecting the development language, identifying the presence of Jupyter notebooks, and locating Python files. This allowed us to filter repositories based on their relevance to our study. The project collection process resulted in 1,900 Python GitHub repositories related to ML services. We randomly selected a statistically significant sample of 320 projects, calculated based on a 95% confidence level and a 5% margin of error for a population of 1,900666https://www.calculator.net/sample-size-calculator.html?type=1&cl=95&ci=5&pp=50&ps=1900&x=Calculate, for further manual analysis. However, we analyzed 57 more projects (leading to a total of 377 projects) to strengthen the reliability and depth of our analysis. Our goal was to ensure broad representation by including repositories related to major cloud providers (i.e., Microsoft, Google, and Amazon) and diverse types of ML services (e.g., text-to-speech, image recognition, and video processing).
    However, we found that not all repositories that mention cloud services or ML-related terms really use ML services. Hence, to maintain the focus on projects that use cloud-based ML services, we excluded repositories that met at least one of the following criteria:
    • Toy or tutorial projects: Repositories created solely for educational purposes (e.g., short workshop demos) were excluded, as they do not represent real-world development practices or deployed systems. For example, a repository [^7] https://github.com/gowhich501/customVisionApidemonstrating how to call an ML service once using a minimal script was excluded even though it mentioned “Azure ML”.
    • Official Toolkits from Cloud Providers: Such repositories [^8] https://github.com/aws-samples/reinvent2021-tsa-bot, [^9] https://github.com/aws-samples/amazon-sagemaker-groundtruth-and-amazon-comprehend-ner-examples, [^10] https://github.com/Azure-Samples/ai-rag-chat-evaluator, often provided by cloud providers (e.g., “AzureML examples” or “aws-sagemaker-notebook-samples”), serve as starter templates and do not reflect actual usage scenarios authored by practitioners. Although they are technically related to ML services, their primary purpose is to present product features rather than complete projects with context-specific implementations.
    • Misleading mentions of ML services without actual use: Some repositories mention cloud providers or ML concepts in their ReadMe files or project descriptions, but do not make use of any cloud-based ML service in their source code. For example, we excluded a project that deployed a microservice on Google Cloud Platform and mentioned “Google AI” in its documentation but did not interact with any ML services (e.g., Vertex AI) in the actual codebase.
    Two of the co-authors (with significant experience in machine learning and software engineering) played the role of evaluators and manually applied the above exclusion criteria independently, yielding a set of 87 ML service-based projects for further analysis. Their inter-rater agreement for this filtering stage, measured using Cohen’s kappa, was 100%, indicating a high level of consistency.
@@
    Projects Analyses: The same co-authors manually and independently analyzed the projects to detect the presence of different ML service misuses. We used Understand [^11] https://scitools.com, a static code analysis tool, to perform a thorough code review and control flow graph generation in complex repositories containing numerous directories. This tool provides different views of the architecture, thus enabling a comprehensive understanding and analysis of the systems.
@@
    Misuses Isolation: As part of the detection process, the two co-authors independently and systematically identified and highlighted instances of ML service misuse, categorizing them by their stage of the ML development pipeline. Each detected misuse is accompanied by a detailed explanation, which provides reasons as to why the identified instance is problematic and a link to the exact line of code within the repository, facilitating quick access for further inspection. For example, to detect “Non specification of early stopping criteria” misuse, the two co-authors independently analyzed the ML service-based system, identified the ML service cloud provider used, verified the training component of the code, and reviewed the recommendations for specifying early stopping criteria based on the official documentation of the cloud provider. If the training component did not adhere to best practices for cost-efficiency and resource optimization, a misuse was detected.
@@
    Detection Agreement: The two co-authors tried to resolve any disagreement during the process through discussion meetings. In cases where a consensus could not be reached, a third coauthor was involved to provide additional insights and resolve disagreements. To assess the level of agreement between the evaluators, we calculated Cohen’s kappa coefficient (banerjee1999beyond, ). We achieved Cohen’s kappa of 84.7%, indicating a high level of agreement among the evaluators and highlighting the reliability of the manual identification process of the occurrences of ML service misuses.
4.4. Consolidation of the Catalog of ML Service Misuses (RQ1)
@@
We curated a preliminary catalog of ML service misuses by combining insights from the research literature, the gray literature, and real-world examples from open-source projects (RQ1). The identification process involved iterative refinement and pre-validation of the identified misuses from each dimension of our multi-vocal study. The two co-authors collaboratively reviewed and evaluated in regular meetings the relevance of the identified misuses, applying the following inclusion criteria: (1) misuses specific to ML service–based systems, excluding those applicable to general service systems; (2) misuses applicable to all cloud providers; (3) misuses with occurrences identified within GitHub projects; and (4) misuses that are potentially detectable through source code analysis. Consensus was reached through a consultation with the third co-author. We also categorized the misuses according to their stage within the ML development pipeline. The final curated set served as the basis for a practitioner survey in which we asked developers about the relevance and occurrence of these misuses in their experience.
4.5. Online Survey with Practitioners (RQ2 & RQ3)
We conducted an online survey with practitioners to validate our catalog and gain better insight into the usage of ML cloud services in the industry (RQ2). This also enabled us to examine and compare the prevalence of ML cloud services misuses across both industry and open-source projects (RQ3). To build the survey, we followed established guidelines in the literature (molleri2020survey, ; fowler2013survey, ) that provide recommendations for different stages of the survey, such as question design, sampling, data collection, and analysis. These recommendations helped us identify and structure the different parts of the survey, craft precise questions, and set suitable response options. The survey consisted of three main phases, explained below.
  • (1)
@@ 4.5. Online Survey with Practitioners (RQ2 & RQ3)
    Preparation of the Online Survey. We created a web-based survey [^12] https://forms.gle/Na3sCzyWt6WmTA7e9 using Google Forms. We designed our survey questions based on the research literature review, the gray literature review, and the analysis of ML service-based systems. Before publishing the survey, we conducted a pilot interview with three professionals and four researchers to identify any duplicated, inconsistent, or unclear questions in the survey before distribution. This also helped to assess the survey length, which might discourage participants from responding, and to estimate the approximate time required to complete the survey. The seven participants went through the survey and suggested minor modifications. To facilitate participants’ understanding of ML service misuses, we included a link to our online catalog [^13] https://ml-service-misuses.GitHub.io/Catalog-Proposal in the survey to further explain each misuse. The final survey consisted of seven sections, primarily about the demographics of the participants, the frequency of usage of ML services, the introduction of ML misuses, the agreement level with the catalog, and the possible reasons for misuses.
  • (2)
@@
    Selection of Participants. We targeted practitioners with experience in ML cloud services. Identifying and inviting such practitioners was challenging due to the key requirement for familiarity with both ML and cloud computing, as well as practical experience with ML services in their work. To select the participants, we relied on (1) information about companies that offer ML cloud services and their clients/partners, (2) the authors’ networks (e.g., former colleagues), and (3) search queries on LinkedIn profiles (e.g., AI Engineer OR Cloud Architect OR Azure certified OR Cloud for AI). For those who were identified as potential practitioners, we sent them invitations via EMail and LinkedIn messages. We chose to limit our invitations to a maximum of five practitioners per organization to ensure a wide representation and mitigate the risk of bias. We invited a total of 421 practitioners to complete the survey and encouraged them to share our invitations within their network. The survey was completed by 50 practitioners, which corresponds to a 12% completion rate, which is somewhat reasonable for practitioner-oriented surveys in software engineering research, as achieving high response rates is challenging in this kind of studies (ghazi2018survey, ).
  • (3)
    Data Validation. To ensure the reliability and accuracy of the survey responses, we employed a multi-step validation process. First, we performed a preliminary screening of the responses to identify incomplete or inconsistent answers. Responses that contained missing critical information or exhibited contradictions were flagged for further verification. To ensure that participants had experience with ML cloud services, we verified their eligibility based on multiple self-reported indicators collected in the survey, including the number of ML-service based projects they had worked on, the specific cloud-based ML technologies used (e.g., SDKs, APIs, and UI), the frequency of using ML services, and their years of experience in AI, software development, and cloud computing. Participants whose responses revealed no relevant background or whose responses were contradictory would have been excluded from further analysis. However, no participants met these exclusion criteria. After that, we extended invitations for follow-up interviews with participants to confirm their responses, validate key findings, clarify ambiguities, and ensure that the survey data accurately reflects the experiences of the participants. By following these validation steps, we aimed to improve the robustness of our findings and minimize potential biases in the survey results.
Statistical Correlation. We measured Spearman’s correlations (ali2022spearman, ) between (a) participants’ experience levels in AI development and the number of ML service misuses they agreed upon, as well as (b) participants’ experience levels and the frequency of encountering each ML service misuse. We used Spearman’s correlation as it does not assume a specific distribution of the data. The goal was to explore whether more experienced practitioners report being aware of, or exposed to, a higher number of ML service misuses in practice. Understanding such trends helps us to assess whether awareness of misuses increases with expertise or, conversely, if misuses are prevalent regardless of experience level. These insights provide useful implications for targeting awareness tools and educational materials toward specific practitioner groups.
Additional Qualitative Insights: To complement the survey data and provide richer context to our findings, we conducted a qualitative analysis of practitioner-oriented online resources. These included blog posts, technical documentation, and Q&A forums such as Stack Overflow, which we used solely for qualitative analysis rather than for catalog construction and consolidation. We also examined research papers discussing challenges and impacts related to cloud-based ML services. Insights from this analysis helped us confirm the reported misuses, understand their introduction reasons and mitigation strategies, and strengthen the interpretation of our survey results.